# Data-centric Prediction Explanation via Kernelized Stein Discrepancy

### v2

@@ Abstract
Existing example-based prediction explanation methods often bridge test and training data points through the model’s parameters or latent representations. While these methods offer clues to the causes of model predictions, they often exhibit innate shortcomings, such as incurring significant computational overhead or producing coarse-grained explanations. This paper presents a Highly-precise and Data-centric Explanation (HD-Explain) prediction explanation method that exploits properties of Kernelized Stein Discrepancy (KSD). Specifically, the KSD uniquely defines a parameterized kernel function for a trained model that encodes model-dependent data correlation. By leveraging the kernel function, one can identify training samples that provide the best predictive support to a test point efficiently. We conducted thorough analyses and experiments across multiple classification domains, where we show that HD-Explain outperforms existing methods from various aspects, including 1) preciseness (fine-grained explanation), 2) consistency, and 3) computation efficiency, leading to a surprisingly simple, effective, and robust prediction explanation solution.
@@ 2.1 Post-hoc Classifier Explanation by Examples
Post-hoc Classifier Explanation by Examples (a.k.a prototypes, representers) refers to a category of classifier explanation approaches that pick a subset of training data points as prediction explanations without accessing the model training process. Its research history spans from the model-intrinsic approach (Molnar, 2020) to the recent impact-based approach (Li et al., 2020).
@@
Model-inherent approaches (Molnar, 2020) refer to machine learning models that are considered interpretable such as $k$-nearest neighbor (Peterson, 2009) or decision tree; For a given test data point, similar data points on the raw feature space can be efficiently selected as explanations through the inherent decision making mechanism of the self-explanatory machine learning models. In fact, attracted by their inherent explanatory power, multiple well-known works attempted to compile complex black-box models into self-explanatory models for enabling prediction explanation (Frosst & Hinton, 2017), while computationally inefficient.
@@
To unlock the general explanatory power applicable to black-box models, multiple later studies suggested to fall back to statistics-based solutions, looking for prototype samples that represent data points that either are common in the dataset or play critical roles in data distribution. MMD-critic (Kim et al., 2016) and Normative and Comparative explanations (Cai et al., 2019) are the well-known examples in this category. Unfortunately, those approaches are often with strong assumption of good prototypes (as of training data distribution) (Li et al., 2020), making their prediction explanations general to the training dataset rather than a trained model instance.
@@
Recently, influence-based methods have emerged as the prevailing technique in model explanation (Li et al., 2020; Nam et al., 2022; Bae et al., 2022; Park et al., 2023). Influence function (Koh & Liang, 2017), as one of the earliest influence-based solutions, bridges the outcome of a prediction task to training data points by, first, evaluating training data’s influence on the model parameters and, then, estimating how model parameter changes affect prediction. Similarly, Representative Point Selection (RPS) (Yeh et al., 2018) builds such an influence chain by fitting the representation theorem, where the weighted product between the representations of test and training samples comes into play. Concerning the computational overhead of previous work, the later solution TracIn (Pruthi et al., 2020) proposed a simple approximation of the influence function via a first-order Taylor expansion (essentially Neural Tangent Kernel (Jacot et al., 2018)), successfully discarding the inverse Hessian matrix from the influence chain formulation. BoostIn (Brophy et al., 2023) further extends TracIn and is dedicated to interpreting the predictions of gradient-boosted decision trees. RPS-LJE (Sui et al., 2021), on the other hand, alleviated the inconsistent explanation problem of RPS through Local Jacobian Expansion. In the latest publication (Tsai et al., 2023), all of the above methods described in this paragraph are identified as special cases of Generalized Representers but with different chosen kernels.
@@ 2.2 Kernelized Stein Discrepancy
The identity can characterize distribution $p(x)$ such that it is often served to assess the goodness-of-fit (Kubokawa, 2024). The above expression could be further abstracted to use function operator $\mathcal{A}_{p}$ (a.k.a Stein operator) such that
@@ 3.1 KSD between Model and Training Data
Inspired by the previous study on Goodness of Fit (Jitkrittum et al., 2020), to unlock KSD support on predictive models, we propose to set $P_{\theta}(\mathbf{x})\equiv P_{D}(\mathbf{x})$ such that the identical marginal distribution would not contribute to the discrepancy between the joint distributions $P_{D}(\mathbf{x},y)$ and $P_{\theta}(\mathbf{x},y)$. In addition, while original distribution $P(\mathbf{x})$ that generates data could be arbitrary complex distribution (that out of modelling scope of the predictive model), we may simply set data point distribution $P_{D}$ (not original distribution $P$) as an Uniform distribution over data points in the dataset under the particular context, given a data point is sampled from a generated dataset uniformly. Although the relaxation appears hasty, we believe that the relaxation is valid in the prediction explanation context (and probably only valid in this particular context) in terms of measuring correlation between a pair of data points, where all train/test data points’ input are valid observation (e.g. images) rather than random continuous valued sample in the same space form an unknown distribution.
@@ 4 Evaluation and Analysis
Datasets: We consider multiple disease classification tasks where diagnosis explanation is highly desired. We also introduced synthetic and benchmark classification datasets to deliver the main idea without the need for medical background knowledge. Concretely, we use CIFAR-10 ($32\times 32\times 3$), Brain Tumor (Magnetic Resonance Imaging, $128\times 128\times 3$), Ovarian Cancer (Histopathology Images, $128\times 128\times 3$) datasets, and SVHN ($32\times 32\times 3$). More details are listed in the Appendix G.
Baselines: The baseline explainers used in our experiments include Influence Function, Representer Point Selection, and TracIn. While other variants of these baseline explainers exist (Barshan et al., 2020; Sui et al., 2021; Chen et al., 2021), we note they don’t offer fundamental performance improvements over the classic ones. In addition, as Influence Function and TracIn face scalability issues, we limit the influence of parameters to the last layer of the model so that they can work with models that contain a large number of parameters. Our experiments use ResNet-18 as the backbone model architecture (with around $11$ million trainable parameters) for all image datasets (see Appendix I for detail on our hardware setup). Finally, we also introduce an HD-Explain variant (HD-Explain*) to match the last layer setting of other baseline models, even though HD-Explain can scale up to the whole model without computation pressure (see Appendix J for details).
@@
4.1 Qualitative Evaluation
Figure 3 shows three test cases of the CIFAR10 classification task that cover different classification outcomes, including high-confident correct prediction, low-confident correct prediction, and low-confident incorrect prediction. For both correct prediction cases, we are confident that HD-Explain provides a better explanation than others in terms of visually matching test data points e.g. brown frogs in Figure 3 (a) and deer on the grass in Figure 3 (b). In contrast, for the misclassified prediction case (Figure 3 (c)), we note the HD-Explain produces an example that does not even belong to the same class as the predicted one. This reflects a low confidence in model’s prediction for the particular test example and highlight a potential error in the prediction. RPS also shows such inconsistency in explanation, which aligns with its claim (Yeh et al., 2018). The other two baseline methods do not offer such properties and still produce explanations that match the predicted label well. It is hard to justify how those training samples support such prediction visually (since no clear shared pattern is obvious to us). In addition, it is interesting to see that Influence Function and TracIn produce near identical explanations, reflecting their similarity in leveraging the perturbation of model parameters.
@@ 4.1 Qualitative Evaluation
Figure 3: Qualitative evaluation of various example-based explanation methods using CIFAR10. We show three scenarios where the target model makes 1) a highly-confident prediction that matches ground truth label, 2) a low-confident prediction that matches ground truth label, 3) low-confident prediction that does not match ground truth label (which is a bird). For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
@@
Figure 4: Qualitative evaluation of various example-based explanation methods using SVHN. We show three scenarios where the target model makes 1-2) a highly-confident prediction that matches ground truth label, 2) a low-confident prediction that matches ground truth label. The model predictive accuracy is near 100%, such that there is no incorrect prediction observed during our random search. For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
Figure 4 provides additional insights on SVHN dataset. HD-Explain again shows a better explanation for producing training samples that appear similar to the test samples (note for the semantic similarity, these explanations should be referred to a medical practitioner). For instance, the explanation of HD-Explain follows the similar environmental context of test points as shown in Figure 4 (a-b). To verify this observation further, we conducted a quantitative evaluation as described in the next section.
@@ 4.2 Quantitative Evaluation
Figure 5(a) shows the hit rate comparison among candidate methods on the four image classification datasets under Noise Injection data augmentation. The existing methods face significant difficulty in retrieving the ideal explanatory sample ($\leq 10\%$), even with such a simple problem setup; only HD-Explain (and its variant) produces a reasonable successful rate ($>80\%$). We further investigate the diversity of explanations across a testset using the Coverage metric. Here, diversity indirectly reflects the granularity of an explanation when accumulated over the test dataset. Figure 5(b) shows the Coverage score, the ratio of explanation samples that are unique over many test points. It turns out that existing solutions produce only 10% - 50% coverage – many test points receive the same set of explanations, disregarding their unique characteristic. This result is in line with our previous observation in Figure 8, where we saw multiple duplicated explanations (as shown with red corner tags). We further observed that the explanations of baselines are often dominated by the class labels; data points predicted as the same class would receive a similar set of explanations. In contrast, HD-Explain shows substantially higher coverage, generating explanations that considers the unique characteristics of each test point.
@@ Appendix C Additional Derivation of Kernelized Stein Discrepancy
Appendix D Discussion: Intuition on why HD-Explain works
[…]
We examine the effect of each term as follows:
    $[\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}||f_{\theta}(\mathbf{x}_{a})]^{\top}$ $[\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}||f_{\theta}(\mathbf{x}_{b})]$
    $=\underbrace{\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}^{\top}\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}}_{\text{similarity of scores (input gradients)}}+\underbrace{f_{\theta}(\mathbf{x}_{a})^{\top}f_{\theta}(\mathbf{x}_{b})}_{\text{similarity of predictions}},$
Appendix E Discussion: Relaxations of KSD Estimation
@@
Appendix F Discussion: Prediction Explain Quality on Healthcare Dataset
@@
Appendix G Dataset Details
@@
Overian Cancer Histopathology (Private) Image 20,000 $128\times 128\times 3$ 5 Yes No
@@
In this paper, we conducted our experiments on five datasets – two synthetic and three benchmark image classification datasets. As the work concerns the trustworthiness of the machine learning model in high-stakes applications, we also introduced medical diagnosis datasets to provide more insight into the potential benefit the proposed work introduced. To train the target machine learning models, we conducted data augmentations to increase the number of training data samples, including random cropping, rotation, shifting, horizontal flipping, and noise injection. Table 2 summarizes more details about the datasets.
Appendix H Data Debugging
@@
Appendix I Hardware setup
@@
Appendix J HD-Explain*
In the main paper, we briefly mentioned HD-Explain* as an HD-Explain variant to match the last layer setting of other baseline models (e.g. RPS). The HD-Explain* is a simple change of HD-Explain in terms of using data representations (the output of the last non-predictive layer of the neural classifier) rather than the raw features. Specifically, we assume a neural network model $f_{\theta}$ could be decomposed into two components $f_{\theta_{2}}\cdot f_{\theta_{1}}$, where $f_{\theta_{1}}$ is a representation encoder and $f_{\theta_{2}}$ is a linear model for prediction. With this decomposition, we define the KSD kernel function for HD-Explain* as
Appendix K Broader Impact
@@
Appendix L Limitation of proposed method

### v3

@@ Abstract
Existing example-based prediction explanation methods often bridge test and training data points through the model’s parameters or latent representations. While these methods offer clues to the causes of model predictions, they often exhibit innate shortcomings, such as incurring significant computational overhead or producing coarse-grained explanations. This paper presents a Highly-precise and Data-centric Explanation (HD-Explain) prediction explanation method that exploits properties of Kernelized Stein Discrepancy (KSD). Specifically, the KSD uniquely defines a parameterized kernel function for a trained model that encodes model-dependent data correlation. By leveraging the kernel function, one can identify training samples that provide the best predictive support to a test point efficiently. We conducted thorough analyses and experiments across multiple classification domains, where we show that HD-Explain outperforms existing methods from various aspects, including 1) preciseness (fine-grained explanation), 2) consistency, and 3) computation efficiency, leading to a surprisingly simple, effective, and robust prediction explanation solution. Source code is available at https://github.com/MahtabSarvmaili/HDExplain.
@@ 2.1 Post-hoc Classifier Explanation by Examples
Post-hoc Classifier Explanation by Examples (a.k.a prototypes, representers) refers to a category of classifier explanation approaches that pick a subset of training data points as prediction explanations without accessing the model training process. Its research history spans from the model-intrinsic approach (see survey  Molnar (2020)) to the recent impact-based approach (Li et al., 2020).
@@
Model-inherent approaches (Molnar, 2020) refer to machine learning models that are considered interpretable such as $k$-nearest neighbor (Peterson, 2009) or decision tree; For a given test data point, similar data points on the raw feature space can be efficiently selected as explanations through the inherent decision making mechanism of the self-explanatory machine learning models. In fact, attracted by their inherent explanatory power, multiple well-known works attempted to compile complex black-box models into self-explanatory models for enabling prediction explanation (Frosst & Hinton, 2017; Yu et al., 2024; Rajani et al., 2020), while computationally inefficient.
@@
To unlock the general explanatory power applicable to black-box models, multiple later studies suggested to fall back to statistics-based solutions, looking for prototype samples that represent data points that either are common in the dataset or play critical roles in data distribution. MMD-critic (Kim et al., 2016) and Normative and Comparative explanations (Cai et al., 2019) are the well-known examples in this category. Unfortunately, those approaches are often with strong assumption of good prototypes (which often overlook the characteristics of trained models) (Li et al., 2020), making their prediction explanations general to the training dataset rather than a trained model instance.
@@
Recently, influence-based methods have emerged as the prevailing technique in model explanation (Li et al., 2020; Nam et al., 2022; Bae et al., 2022; Park et al., 2023a). Influence function (Koh & Liang, 2017), as one of the earliest influence-based solutions, bridges the outcome of a prediction task to training data points by, first, evaluating training data’s influence on the model parameters and, then, estimating how model parameter changes affect prediction. Similarly, Representative Point Selection (RPS) (Yeh et al., 2018) builds such an influence chain by fitting the representation theorem, where the weighted product between the representations of test and training samples comes into play. Concerning the computational overhead of previous work, the later solution TracIn (Pruthi et al., 2020) proposed a simple approximation of the influence function via a first-order Taylor expansion (essentially Neural Tangent Kernel (Jacot et al., 2018)), successfully discarding the inverse Hessian matrix from the influence chain formulation. BoostIn (Brophy et al., 2023) further extends TracIn and is dedicated to interpreting the predictions of gradient-boosted decision trees. RPS-LJE (Sui et al., 2021), on the other hand, alleviated the inconsistent explanation problem of RPS through Local Jacobian Expansion. In the latest publication (Tsai et al., 2023), all of the above methods described in this paragraph are identified as special cases of Generalized Representers but with different chosen kernels.
@@ 2.2 Kernelized Stein Discrepancy
The identity can characterize distribution $p(x)$ such that it is often served to assess the goodness-of-fit  (Kubokawa, 2024) of the model. The above expression could be further abstracted to use function operator $\mathcal{A}_{p}$ (a.k.a Stein operator) such that
@@ 3.1 KSD between Model and Training Data
Inspired by the previous study on Goodness of Fit (Jitkrittum et al., 2020), to unlock KSD support on predictive models, we propose to set $P_{\theta}(\mathbf{x})\equiv P_{D}(\mathbf{x})$ such that the identical marginal distribution would not contribute to the discrepancy between the joint distributions $P_{D}(\mathbf{x},y)$ and $P_{\theta}(\mathbf{x},y)$. In addition, while original distribution $P(\mathbf{x})$ that generates data could be arbitrary complex distribution (that is out of modeling scope of the predictive model), we may simply set data point distribution $P_{D}$ (not original distribution $P$) as an Uniform distribution over data points in the dataset under the particular context, given a data point is sampled from a generated dataset uniformly. Although the relaxation appears hasty, we believe that the relaxation is valid in the prediction explanation context (and probably only valid in this particular context) in terms of measuring correlation between a pair of data points, where all train/test data points’ input are valid observation (e.g. images) rather than random continuous valued sample in the same space form an unknown distribution.
@@ 4 Evaluation and Analysis
Datasets: We consider multiple disease classification tasks where diagnosis explanation is highly desired. We also introduced synthetic and benchmark classification datasets to deliver the main idea without the need for medical background knowledge. Concretely, we use CIFAR-10 ($32\times 32\times 3$), Brain Tumor (Magnetic Resonance Imaging, $128\times 128\times 3$), Ovarian Cancer (Histopathology Images, $128\times 128\times 3$) datasets, and SVHN ($32\times 32\times 3$). More details are listed in the Appendix F.
Baselines: The baseline explainers used in our experiments include Influence Function, Representer Point Selection, and TracIn. While other variants of these baseline explainers exist (Barshan et al., 2020; Sui et al., 2021; Chen et al., 2021), we note they don’t offer fundamental performance improvements over the classic ones. In addition, as Influence Function and TracIn face scalability issues, we limit the influence of parameters to the last layer of the model so that they can work with models that contain a large number of parameters. Our experiments use ResNet-18 as the backbone model architecture (with around $11$ million trainable parameters) for all image datasets (see Appendix H for detail on our hardware setup).
Finally, we also introduce an HD-Explain variant (HD-Explain*) to match the last layer setting of other baseline models, even though HD-Explain can scale up to the whole model without computation pressure. The HD-Explain* is a simple change of HD-Explain in terms of using data representations (the output of the last non-predictive layer of the neural classifier) rather than the raw features. Specifically, we assume a neural network model $f_{\theta}$ could be decomposed into two components $f_{\theta_{2}}\cdot f_{\theta_{1}}$, where $f_{\theta_{1}}$ is a representation encoder and $f_{\theta_{2}}$ is a linear model for prediction. With this decomposition, we define the KSD kernel function for HD-Explain* as
@@
Figure 3: Qualitative evaluation of various example-based explanation methods using CIFAR10. We show three scenarios where the target model makes a) a highly-confident prediction that matches ground truth label, b) a low-confident prediction that matches ground truth label, c) low-confident prediction that does not match ground truth label (which is a bird). For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example.
@@
Figure 4: Qualitative evaluation of various example-based explanation methods using SVHN. We show two scenarios where the target model makes a-b) a highly-confident prediction that matches ground truth label, c) a low-confident prediction that matches ground truth label. For each sub plot, we show top-3 influential training data points picked by the explanation methods for the test example. We include two samples of high-confidence correct predictions to show the overlap of explanations.
4.1 Qualitative Evaluation
Figure 3 shows three test cases of the CIFAR10 classification task that cover different classification outcomes, including high-confidence correct prediction, low-confident correct prediction, and low-confident incorrect prediction. For both correct prediction cases, we are confident that HD-Explain provides a better explanation than others in terms of visually matching test data points e.g. brown frogs in Figure 3 (a) and deer on the grass in Figure 3 (b). In contrast, for the misclassified prediction case (as shown in Figure 3 (c)), we note the HD-Explain produces an example that does not even belong to the same class as the predicted one. pecifically, the predicted class is cat (while the ground truth label is bird), and HD-Explain generates an explanation sample from the deer class. This reflects a low confidence in model’s prediction for the particular test example and highlight a potential error in the prediction. RPS also shows such inconsistency in explanation, which aligns with its claim (Yeh et al., 2018). The other two baseline methods do not offer such properties and still produce explanations that match the predicted label well. It is hard to justify how those training samples support such prediction visually (since no clear shared pattern is obvious to us). In addition, it is interesting to see that Influence Function and TracIn produce near identical explanations, reflecting their similarity in leveraging the perturbation of model parameters.
Figure 4 provides additional insights on SVHN dataset. HD-Explain again shows a better explanation for producing training samples that appear similar to the test samples. In addition, we notice that RPS produces the same set of explanations for different sample cases, as shown in Fig. 4 (a-b), which reveals its limitation in providing instance-level explanations. To verify this observation further, we conducted a quantitative evaluation as described in the next section.
@@ 4.2 Quantitative Evaluation
Figure 5(a) shows the hit rate comparison among candidate methods on the four image classification datasets under Noise Injection data augmentation. The existing methods face significant difficulty in retrieving the ideal explanatory sample ($\leq 10\%$), even with such a simple problem setup; only HD-Explain (and its variant) produces a reasonable successful rate ($>80\%$). We further investigate the diversity of explanations across a testset using the Coverage metric. Here, diversity indirectly reflects the granularity of an explanation when accumulated over the test dataset. Figure 5(b) shows the Coverage score, the ratio of explanation samples that are unique over many test points. It turns out that existing solutions produce only 10% - 50% coverage – many test points receive the same set of explanations, disregarding their unique characteristic. We further observed that the explanations of baselines are often dominated by the class labels; data points predicted as the same class would receive a similar set of explanations. In contrast, HD-Explain shows substantially higher coverage, generating explanations that considers the unique characteristics of each test point.
@@ 4.3 Kernel Options
4.4 Discussion: Intuition on why HD-Explain works
After showing HD-Explain’s empirical performance, we now present our understanding of how HD-Explain finds the explanations for the prediction of a test data point. In particular, we want to understand why the approach is faithful to the pre-trained model.
[…]
We examine the effect of each term as follows:
    $[\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}$ $||\log f_{\theta}(\mathbf{x}_{a})]^{\top}[\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}||\log f_{\theta}(\mathbf{x}_{b})]$
    $=\underbrace{\nabla_{\mathbf{x}_{a}}\log f_{\theta}(\mathbf{x}_{a})_{y_{a}}^{\top}\nabla_{\mathbf{x}_{b}}\log f_{\theta}(\mathbf{x}_{b})_{y_{b}}}_{\text{similarity of scores (input gradients)}}+\underbrace{\log f_{\theta}(\mathbf{x}_{a})^{\top}\log f_{\theta}(\mathbf{x}_{b})}_{\text{similarity of predictions}},$
@@ 5 Conclusion
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Foundation of Innovation (CFI), and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
@@ Appendix C Additional Derivation of Kernelized Stein Discrepancy
Appendix D Discussion: Relaxations of KSD Estimation
@@
Appendix E Discussion: Prediction Explain Quality on Healthcare Dataset
@@
Appendix F Dataset Details
@@
Overian Cancer Histopathology Image 20,000 $128\times 128\times 3$ 5 Yes No
@@
In this paper, we conducted our experiments on five datasets – two synthetic and three benchmark image classification datasets. As the work concerns the trustworthiness of the machine learning model in high-stakes applications, we also introduced medical diagnosis datasets (Bashashati et al., 2023) to provide more insight into the potential benefit the proposed work introduced. To train the target machine learning models, we conducted data augmentations to increase the number of training data samples, including random cropping, rotation, shifting, horizontal flipping, and noise injection. Table 2 summarizes more details about the datasets.
Appendix G Data Debugging
@@
Appendix H Hardware setup
@@
Appendix I Broader Impact
@@
Appendix J Limitation of proposed method
@@
In the main paper, we evaluated HD-Explain with three well-known kernel options, including Linear, RBF, and Inverse Multi-Quadric (IMQ). The purpose was to demonstrate the impact of kernel choices on the performance of HD-Explain. While we show that IMQ performs best in our experimental environment, we want to highlight that the selection of the kernel may influence the HD-Explain’s performance in practice. Hence, one may need to conduct empirical analysis on which kernel to use before applying HD-Explain in production.
Appendix K Relation to Data Attribution Estimation
Data attribution estimation is closely related to the sample-based prediction explanation but a different concept. As stated by Park et al. (2023b), the definition of data attribution is as follows:
Definition K.1 (Data attribution).
Consider an ordered training set of examples $S=\{z_{1},\ldots,z_{n}\}$ and a model output function $f({z};{\theta})$. A data attribution method $\tau(z,S)$ is a function $\tau:\mathcal{Z}\times\mathcal{Z}^{n}\to\mathbb{R}^{n}$ that, for any example $z\in\mathcal{Z}$ and a training set $S$, assigns a (real-valued) score to each training input $z_{i}\in S$ indicating its importance to the model output $f(z;\theta^{*}(S))$. When the second argument $S$ is clear from the context, we will omit the second argument and simply write $\tau(z)$.
In particular, data attribution estimation faces a more generalized problem that does not explain the importance of data for a specific pre-trained model but the importance of a family of models of the same architecture or function. The star ($\theta^{*}$) refers to the potentially optimal model that can be trained on the dataset $S$. Indeed, if we examine the two approaches in the data attribution estimation literature (data modelling (Ilyas et al., 2022) and TRAK (Park et al., 2023b)), we note both approaches require either training multiple models on a subset of data points or introducing various aggressive approximations, such as (1) linear Taylor expansion, 2) random projection, and 3) newton approximation. Both data and model manipulations will cause unfaithfulness to the pre-trained model. In fact, the data modelling algorithm (Ilyas et al., 2022) does not involve any pre-trained model but directly optimizes the linear data modelling score as follows.
$\tau_{\textsc{dm}}(z)\coloneqq\min_{\beta\in\mathbb{R}^{n}}\frac{1}{m}\sum_{i=1}^{m}({{\beta}^{\top}1_{S_{i}}-f(z;{\theta^{*}(S_{i})})}^{2}+\lambda\|\beta\|_{1}.$
Appendix L HD-Explain: Explanation Process
The following algorithm shows the algorithm of HD-Explain in pseudocode.
Algorithm 1 HD-Explain
1:Training set $D$, Test input $\mathbf{x}_{\text{test}}$, and classifier model $f_{\theta}$
2:Sample based explanations $D_{\text{explain}}\subset D$
3:
4:Step 1: Caching $\triangleright$ Reduce redundant computation
5:initialize empty list $c\leftarrow[]$
6:for $(\mathbf{x}_{i},y_{i})\in D$ do
7:     $\mathbf{p}_{i}\leftarrow f_{\theta}(\mathbf{x}_{i})$
8:     $\mathbf{g}_{i}\leftarrow\nabla_{\mathbf{x}_{i}}\log f_{\theta}(\mathbf{x}_{i})_{y_{i}}$
9:     c.add($[\mathbf{x}_{i},y_{i},\mathbf{p}_{i},\mathbf{g}_{i}]$)
10:end for
11:
12:Step 2: Prediction Contribution of Each Training Data
13:Given test input $\mathbf{x}_{\text{test}}$
14:$\mathbf{p}_{\text{test}}\leftarrow f_{\theta}(\mathbf{x}_{\text{test}})$
15:$\hat{y}_{\text{test}}\leftarrow\arg\!\max\mathbf{p}_{\text{test}}$triangleright Best Predicted Label
16:$\mathbf{g}_{\text{test}}\leftarrow\nabla_{\mathbf{x}_{\text{test}}}\log f_{\theta}(\mathbf{x}_{\text{test}})_{\hat{y}_{\text{test}}}$
17:c.add($[\mathbf{x}_{\text{test}},y_{\text{test}},\mathbf{p}_{\text{test}},\mathbf{g}_{\text{test}}]$) $\nabla_{\mathbf{x}_{\text{test}},\hat{\mathbf{y}}_{\text{test}}}\log P_{\theta}(\mathbf{x}_{i},\hat{\mathbf{y}}_{i})\leftarrow[\mathbf{g}_{\text{test}}||\mathbf{p}_{\text{test}}]$
18:for $(\mathbf{x}_{i},y_{i})\in D$ do
19:     $\nabla_{\mathbf{x}_{i},\mathbf{y}_{i}}\log P_{\theta}(\mathbf{x}_{i},\mathbf{y}_{i})\leftarrow[\mathbf{g}_{i}||\mathbf{p}_{i}]$triangleright Cache-able if needed
20:     $\kappa_{\theta}((\mathbf{x}_{i},y_{i}),(\mathbf{x}_{\text{test}},\hat{y}_{\text{test}}))\leftarrow$ Equation 1
21:end for
22:$D_{\text{explain}}\leftarrow\text{argsort}([\kappa_{\theta}((\mathbf{x}_{i},y_{i}),(\mathbf{x}_{\text{test}},\hat{y}_{\text{test}}))\text{ for }i\in|D|])$
Appendix M Additional Qualitative Evaluation Examples for SVHN
(a) Low-Conf Incorrect Predict
Figure 11: Qualitative evaluation of various example-based explanation methods using SVHN. We show the missing scenario in the main paper, where the target model makes low-confident prediction that does not match ground truth label (which is a $7$). For the sub plot, we show top-3 influential training data points picked by the explanation methods for the test example. The observation is similar to that of CIFAR-10.